OpenStudy (anonymous):

Find all values for x that must be excluded from the domain of the expression: 2x(x + 1)/3x2 + 5x + 2

OpenStudy (anonymous):

Oops, wrote it wrong.

OpenStudy (anonymous):

Did you mean? \[\huge \frac{\frac{2x}{x + 1}}{3x^2} + 5x + 2\] The obvious part here is that no denominator must = 0, but first I would simplify.

OpenStudy (anonymous):

You will have to exclude anything that causes a "divide by zero" error when you have this in simplified form. :-) Right now those two things are x+1=0 x=-1 and 3x^2 = 0 x^2=0 x=0

OpenStudy (anonymous):

Thanks!